# Interplay of Zeeman field, Rashba spin-orbit interaction, and superconductivity: Transition temperature and quasiparticle excitations

### v1

@@ IV.1 Opposite-spin pairing state
$\mathbf{d}(\mathbf{k})=\Delta\sin\theta_{\mathbf{k}}e^{i\varphi_{\mathbf{k}}}\hat{z},$ (32a)
@@ IV.1.1 Superconducting transition temperature
$\frac{T_{c}}{T_{c0}}=\left[\frac{e^{-\gamma}}{4}\cdot\frac{2\pi k_{B}T_{c0}}{gk_{F}}\right]^{\frac{2}{3}},$ (34)
under the condition $k_{B}T_{c}\ll gk_{F}/2\pi$. Refer to Appendix H for a comprehensive set of calculations. On the other hand, in the case of the pairing state $(k_{x}+ik_{y})\hat{z}$, this continuous phase transition turns into a first-order transition when $gk_{F}\gtrsim{}1.23{}k_{B}T_{c0}$.
@@ IV.1.2 Quasiparticle excitation energy
Here, $d_{z}(\mathbf{k})=\Delta\sin\theta_{\mathbf{k}}e^{i\varphi_{\mathbf{k}}}$ corresponds to the pairing state $(k_{x}+ik_{y})\hat{z}$, while $d_{z}(\mathbf{k})=\Delta\cos\theta_{\mathbf{k}}$ describes the $k_{z}\hat{z}$ pairing state, as outlined in Eq. (32). From Eqs. (36) and (37), it can be seen that:
@@ IV.2.1 Superconducting transition temperature
$\frac{T_{c}}{T_{c0}}=\sqrt{3}e^{-\frac{5}{6}}\cdot\frac{H_{0}}{H_{\perp}}.$ (44)
For detailed calculations, refer to Appendix H.
@@
$\frac{T_{c}}{T_{c0}}=\left[\frac{\pi e^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{gk_{F}}\right]^{\frac{1}{4}}$ (45)
@@
$\frac{T_{c}}{T_{c0}}=\left[\frac{\pi e^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{gk_{F}}\right]^{\frac{3}{2}}$ (46)
@@ V Summary and Discussions
  • Orbital effect of a magnetic field: In addition to the Zeeman effect considered in this paper, which concerns the electron spin degrees of freedom, a magnetic field also induces a diamagnetic current known as the orbital effect. In a type-I superconductor, a weak external magnetic field results in the complete expulsion of the magnetic field from the bulk. However, our theory is relevant for the intermediate state in type-I superconductors in a strong magnetic field and applies to type-II superconductors when $H_{c1}<H<H_{c2}$ [35, 1]. Moreover, our theory can also be applied to systems where ferromagnetism and superconductivity coexist [36] or hybridize [37] at the microscopic level.
@@ Appendix B Intraband and interband pairing functions
$\begin{split}&\Delta_{++}=-\Delta_{--}=-\Delta{}e^{-i\varphi_{\mathbf{k}}}\sin{\frac{\Theta_{\mathbf{k}}+\Theta_{-\mathbf{k}}}{2}},\\ &\Delta_{+-}=\Delta_{-+}=\Delta{}e^{-i\varphi_{\mathbf{k}}}\cos{\frac{\Theta_{\mathbf{k}}+\Theta_{-\mathbf{k}}}{2}}.\end{split}$
@@
Appendix C Bogoliubov transformation: $s$-wave pairing state when only Zeeman field or Rashba SOC is present
C.1 In a Zeeman field
@@ C.2 In the presence of Rashba SOC
$u_{\mathbf{k}}=\left\{\begin{aligned} &\begin{pmatrix}a_{\mathbf{k}+}\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ a_{\mathbf{k}+}\sin{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}\end{pmatrix},\qquad&\mathbf{k}\in I_{1}\\ &\begin{pmatrix}a_{\mathbf{k}+}\cos{\frac{\theta_{\mathbf{k}}}{2}}&a_{\mathbf{k}-}\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ a_{\mathbf{k}+}\sin{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}&-a_{\mathbf{k}-}\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}\end{pmatrix},&\mathbf{k}\in I_{2}\\ &\begin{pmatrix}\cos{\frac{\theta_{\mathbf{k}}}{2}}&a_{\mathbf{k}-}\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ \sin{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}&-a_{\mathbf{k}-}\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{i\varphi_{\mathbf{k}}}\end{pmatrix},&\mathbf{k}\in I_{3}\end{aligned}\right.,\quad v_{\mathbf{k}}=\left\{\begin{aligned} &\begin{pmatrix}b_{\mathbf{k}+}\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{-i\varphi_{\mathbf{k}}}&0\\ b_{\mathbf{k}+}\sin{\frac{\theta_{\mathbf{k}}}{2}}&0\end{pmatrix},\quad&\mathbf{k}\in I_{1}\\ &\begin{pmatrix}b_{\mathbf{k}+}\cos{\frac{\theta_{\mathbf{k}}}{2}}e^{-i\varphi_{\mathbf{k}}}&-b_{\mathbf{k}-}\sin{\frac{\theta_{\mathbf{k}}}{2}}e^{-i\varphi_{\mathbf{k}}}\\ b_{\mathbf{k}+}\sin{\frac{\theta_{\mathbf{k}}}{2}}&b_{\mathbf{k}-}\cos{\frac{\theta_{\mathbf{k}}}{2}}\end{pmatrix},&\mathbf{k}\in I_{2}\\ &\begin{pmatrix}0&-b_{\mathbf{k}-}\sin{\frac{\theta_{\mathbf{k}}}{2}}e^{-i\varphi_{\mathbf{k}}}\\ 0&b_{\mathbf{k}-}\cos{\frac{\theta_{\mathbf{k}}}{2}}\end{pmatrix},\qquad&\mathbf{k}\in I_{3}\end{aligned}\right.,$ (66)
@@
Appendix D Finite total momentum $s$-wave pairing states
@@
$F_{s}=-\frac{1}{\beta}\sum_{\mathbf{k}}\left[\ln(1+e^{-\beta E_{\mathbf{k}+\frac{\mathbf{q}}{2},+}})+\ln(1+e^{-\beta E_{\mathbf{k}+\frac{\mathbf{q}}{2},-}})\right]+\sum_{\mathbf{k}}\left[\xi_{-\mathbf{k}+\frac{\mathbf{q}}{2}}-\frac{1}{2}(E_{-\mathbf{k}+\frac{\mathbf{q}}{2},+}+E_{-\mathbf{k}+\frac{\mathbf{q}}{2},-})\right]+\frac{|\Delta_{\mathbf{q}}|^{2}}{V}.$ (72)
@@
Appendix E Spin quantization axes and pairing functions
@@
Appendix F Bogoliubov transformation: $p$-wave pairing states in a single Zeeman field
@@
$H_{\mathbf{k}}=\begin{pmatrix}\xi_{\mathbf{k}}+\mu_{B}|\mathbf{H}|\cos\theta&\mu_{B}|\mathbf{H}|\sin\theta e^{-i\phi}&[-d^{\prime}_{x}(\mathbf{k})+id^{\prime}_{y}(\mathbf{k})]e^{i\eta_{\mathbf{k}}}&d^{\prime}_{z}(\mathbf{k})e^{i\eta_{\mathbf{k}}}\\ \mu_{B}|\mathbf{H}|\sin\theta e^{i\phi}&\xi_{\mathbf{k}}-\mu_{B}|\mathbf{H}|\cos\theta_{\mathbf{k}}&d^{\prime}_{z}(\mathbf{k})e^{i\eta_{\mathbf{k}}}&[d^{\prime}_{x}(\mathbf{k})+id^{\prime}_{y}(\mathbf{k})]e^{i\eta_{\mathbf{k}}}\\ [-d^{\prime}_{x}(\mathbf{k})-id^{\prime}_{y}(\mathbf{k})]e^{-i\eta_{\mathbf{k}}}&d^{\prime}_{z}(\mathbf{k})e^{-i\eta_{\mathbf{k}}}&-\xi_{\mathbf{k}}-\mu_{B}|\mathbf{H}|\cos\theta&-\mu_{B}|\mathbf{H}|\sin\theta_{\mathbf{k}}e^{i\phi}\\ d^{\prime}_{z}(\mathbf{k})e^{-i\eta_{\mathbf{k}}}&[d^{\prime}_{x}(\mathbf{k})-id^{\prime}_{y}(\mathbf{k})]e^{-i\eta_{\mathbf{k}}}&-\mu_{B}|\mathbf{H}|\sin\theta e^{-i\phi}&-\xi_{\mathbf{k}}+\mu_{B}|\mathbf{H}|\cos\theta\end{pmatrix},$ (78)
@@
$\begin{split}\mathbf{d}(\mathbf{k})&=\mathbf{d}^{\prime}(\mathbf{k})e^{i\eta_{\mathbf{k}}},\\ d_{x}(\mathbf{k})&=d_{x}^{\prime}(\mathbf{k})e^{i\eta_{\mathbf{k}}},\\ d_{y}(\mathbf{k})&=d_{y}^{\prime}(\mathbf{k})e^{i\eta_{\mathbf{k}}},\\ d_{z}(\mathbf{k})&=d_{z}^{\prime}(\mathbf{k})e^{i\eta_{\mathbf{k}}}.\end{split}$ (79)
@@
$=$ $\begin{pmatrix}\cos{\frac{\theta}{2}}&\sin{\frac{\theta}{2}}&0&0\\ \sin{\frac{\theta}{2}}e^{i\phi}&-\cos{\frac{\theta}{2}}e^{i\phi}&0&0\\ 0&0&\cos{\frac{\theta}{2}}&\sin{\frac{\theta}{2}}\\ 0&0&\sin{\frac{\theta}{2}}e^{-i\phi}&-\cos{\frac{\theta}{2}}e^{-i\phi}\end{pmatrix}\begin{pmatrix}c_{\mathbf{k}+}\\ c_{\mathbf{k}-}\\ c_{-\mathbf{k}+}^{\dagger}\\ c_{-\mathbf{k}-}^{\dagger}\end{pmatrix},$
@@
$H^{\prime}_{\mathbf{k}}=\begin{pmatrix}\xi_{\mathbf{k}}+\mu_{B}|\mathbf{H}|&0&[-d_{1}(\mathbf{k})+id_{2}(\mathbf{k})]e^{i\gamma_{\mathbf{k}}}&d_{3}(\mathbf{k})e^{i\gamma_{\mathbf{k}}}\\ 0&\xi_{\mathbf{k}}-\mu_{B}|\mathbf{H}|&d_{3}(\mathbf{k})e^{i\gamma_{\mathbf{k}}}&[d_{1}(\mathbf{k})+id_{2}(\mathbf{k})]e^{i\gamma_{\mathbf{k}}}\\ [-d_{1}(\mathbf{k})-id_{2}(\mathbf{k})]e^{-i\gamma_{\mathbf{k}}}&d_{3}(\mathbf{k})e^{-i\gamma_{\mathbf{k}}}&-\xi_{\mathbf{k}}-\mu_{B}|\mathbf{H}|&0\\ d_{3}(\mathbf{k})e^{-i\gamma_{\mathbf{k}}}&[d_{1}(\mathbf{k})-id_{2}(\mathbf{k})]e^{-i\gamma_{\mathbf{k}}}&0&-\xi_{\mathbf{k}}+\mu_{B}|\mathbf{H}|\end{pmatrix},$ (81)

### v2

@@ IV.1 Opposite-spin pairing state
$\mathbf{d}(\mathbf{k})=\Delta\sin\theta_{\mathbf{k}}\mathrm{e}^{i\varphi_{\mathbf{k}}}\hat{z},$ (29a)
@@ IV.1.1 Superconducting transition temperature
$\frac{T_{c}}{T_{c0}}=\left[\frac{\mathrm{e}^{-\gamma}}{4}\cdot\frac{2\pi k_{B}T_{c0}}{gk_{F}}\right]^{\frac{2}{3}},$ (31)
under the condition $k_{B}T_{c}\ll gk_{F}/2\pi$. Refer to Appendix I for a comprehensive set of calculations. On the other hand, in the case of the pairing state $(k_{x}+ik_{y})\hat{z}$, this continuous phase transition turns into a first-order transition when $gk_{F}\gtrsim{}1.23{}k_{B}T_{c0}$.
@@ IV.1.2 Quasiparticle excitation energy
Here, $d_{z}(\mathbf{k})=\Delta\sin\theta_{\mathbf{k}}\mathrm{e}^{i\varphi_{\mathbf{k}}}$ corresponds to the pairing state $(k_{x}+ik_{y})\hat{z}$, while $d_{z}(\mathbf{k})=\Delta\cos\theta_{\mathbf{k}}$ describes the $k_{z}\hat{z}$ pairing state, as outlined in Eq. (29). From Eqs. (33) and (34), it can be seen that:
@@ IV.2.1 Superconducting transition temperature
$\frac{T_{c}}{T_{c0}}=\sqrt{3}\mathrm{e}^{-\frac{5}{6}}\cdot\frac{H_{0}}{H_{\perp}}.$ (41)
For detailed calculations, refer to Appendix I.
@@
$\frac{T_{c}}{T_{c0}}=\left[\frac{\pi\mathrm{e}^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{gk_{F}}\right]^{\frac{1}{4}}$ (42)
@@
$\frac{T_{c}}{T_{c0}}=\left[\frac{\pi\mathrm{e}^{-\gamma}}{2}\cdot\frac{k_{B}T_{c0}}{gk_{F}}\right]^{\frac{3}{2}}$ (43)
@@ V Summary and Discussions
  • Feasibility of a uniform Zeeman field: In addition to the Zeeman effect considered in this paper, which concerns the electron spin degrees of freedom, a magnetic field also induces a diamagnetic current known as the orbital effect. In a type-I superconductor, a weak external magnetic field results in the complete expulsion of the magnetic field from the bulk. However, our theory is relevant for the intermediate state in type-I superconductors in a strong magnetic field and applies to type-II superconductors when $H_{c1}<H<H_{c2}$ [35, 1]. Moreover, our theory can also be applied to systems where ferromagnetism and superconductivity coexist [36] or hybridize [37] at the microscopic level. In such systems, it is possible to expand the calculations for the Zeeman field to encompass local magnetization.
@@ Appendix B Intraband and interband pairing functions
$\begin{split}&\Delta_{++}=-\Delta_{--}=-\Delta{}\mathrm{e}^{-i\varphi_{\mathbf{k}}}\sin{\frac{\Theta_{\mathbf{k}}+\Theta_{-\mathbf{k}}}{2}},\\ &\Delta_{+-}=\Delta_{-+}=\Delta{}\mathrm{e}^{-i\varphi_{\mathbf{k}}}\cos{\frac{\Theta_{\mathbf{k}}+\Theta_{-\mathbf{k}}}{2}}.\end{split}$
@@
Appendix C Gap equation and free energy: uniform $s$-wave pairing state
For the $s$-wave pairing function specified in Eq. (11), the gap equation (7) simplifies to the form
$F_{s}-F_{n}=-\frac{1}{\beta}\sum_{\mathbf{k}}\ln\frac{(1+\mathrm{e}^{-\beta E_{\mathbf{k}+}})(1+\mathrm{e}^{-\beta E_{\mathbf{k}-}})}{(1+\mathrm{e}^{-\beta\xi_{\mathbf{k}+}})(1+\mathrm{e}^{-\beta\xi_{\mathbf{k}-}})}+\frac{\lvert\Delta\lvert^{2}}{V_{0}}+\sum_{\mathbf{k}}\left[\xi_{\mathbf{k}}-\frac{E_{-\mathbf{k}+}+E_{-\mathbf{k}-}}{2}\right],$ (61)
$F_{s}(T=T_{c})-F_{n}(T=T_{c})=0,$ (62)
where $F_{s}-F_{n}$ is specified in Eq. (61).
Appendix D Bogoliubov transformation: $s$-wave pairing state when only Zeeman field or Rashba SOC is present
D.1 In a Zeeman field
@@ D.2 In the presence of Rashba SOC
$u_{\mathbf{k}}=\left\{\begin{aligned} &\begin{pmatrix}a_{\mathbf{k}+}\cos{\frac{\theta_{\mathbf{k}}}{2}}&\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ a_{\mathbf{k}+}\sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}&-\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}\end{pmatrix},\qquad&\mathbf{k}\in I_{1}\\ &\begin{pmatrix}a_{\mathbf{k}+}\cos{\frac{\theta_{\mathbf{k}}}{2}}&a_{\mathbf{k}-}\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ a_{\mathbf{k}+}\sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}&-a_{\mathbf{k}-}\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}\end{pmatrix},&\mathbf{k}\in I_{2}\\ &\begin{pmatrix}\cos{\frac{\theta_{\mathbf{k}}}{2}}&a_{\mathbf{k}-}\sin{\frac{\theta_{\mathbf{k}}}{2}}\\ \sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}&-a_{\mathbf{k}-}\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{i\varphi_{\mathbf{k}}}\end{pmatrix},&\mathbf{k}\in I_{3}\end{aligned}\right.,\quad v_{\mathbf{k}}=\left\{\begin{aligned} &\begin{pmatrix}b_{\mathbf{k}+}\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}&0\\ b_{\mathbf{k}+}\sin{\frac{\theta_{\mathbf{k}}}{2}}&0\end{pmatrix},\quad&\mathbf{k}\in I_{1}\\ &\begin{pmatrix}b_{\mathbf{k}+}\cos{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}&-b_{\mathbf{k}-}\sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}\\ b_{\mathbf{k}+}\sin{\frac{\theta_{\mathbf{k}}}{2}}&b_{\mathbf{k}-}\cos{\frac{\theta_{\mathbf{k}}}{2}}\end{pmatrix},&\mathbf{k}\in I_{2}\\ &\begin{pmatrix}0&-b_{\mathbf{k}-}\sin{\frac{\theta_{\mathbf{k}}}{2}}\mathrm{e}^{-i\varphi_{\mathbf{k}}}\\ 0&b_{\mathbf{k}-}\cos{\frac{\theta_{\mathbf{k}}}{2}}\end{pmatrix},\qquad&\mathbf{k}\in I_{3}\end{aligned}\right.,$ (66)
@@
Appendix E Finite total momentum $s$-wave pairing states
@@
$F_{s}=-\frac{1}{\beta}\sum_{\mathbf{k}}\left[\ln(1+\mathrm{e}^{-\beta E_{\mathbf{k}+\frac{\mathbf{q}}{2},+}})+\ln(1+\mathrm{e}^{-\beta E_{\mathbf{k}+\frac{\mathbf{q}}{2},-}})\right]+\sum_{\mathbf{k}}\left[\xi_{-\mathbf{k}+\frac{\mathbf{q}}{2}}-\frac{1}{2}(E_{-\mathbf{k}+\frac{\mathbf{q}}{2},+}+E_{-\mathbf{k}+\frac{\mathbf{q}}{2},-})\right]+\frac{|\Delta_{\mathbf{q}}|^{2}}{V}.$ (72)
@@
Appendix F Spin quantization axes and pairing functions
@@
Appendix G Bogoliubov transformation: $p$-wave pairing states in a single Zeeman field
@@
$H_{\mathbf{k}}=\begin{pmatrix}\xi_{\mathbf{k}}+\mu_{B}|\mathbf{H}|\cos\theta&\mu_{B}|\mathbf{H}|\sin\theta\mathrm{e}^{-i\phi}&[-d^{\prime}_{x}(\mathbf{k})+id^{\prime}_{y}(\mathbf{k})]\mathrm{e}^{i\eta_{\mathbf{k}}}&d^{\prime}_{z}(\mathbf{k})\mathrm{e}^{i\eta_{\mathbf{k}}}\\ \mu_{B}|\mathbf{H}|\sin\theta\mathrm{e}^{i\phi}&\xi_{\mathbf{k}}-\mu_{B}|\mathbf{H}|\cos\theta_{\mathbf{k}}&d^{\prime}_{z}(\mathbf{k})\mathrm{e}^{i\eta_{\mathbf{k}}}&[d^{\prime}_{x}(\mathbf{k})+id^{\prime}_{y}(\mathbf{k})]\mathrm{e}^{i\eta_{\mathbf{k}}}\\ [-d^{\prime}_{x}(\mathbf{k})-id^{\prime}_{y}(\mathbf{k})]\mathrm{e}^{-i\eta_{\mathbf{k}}}&d^{\prime}_{z}(\mathbf{k})\mathrm{e}^{-i\eta_{\mathbf{k}}}&-\xi_{\mathbf{k}}-\mu_{B}|\mathbf{H}|\cos\theta&-\mu_{B}|\mathbf{H}|\sin\theta_{\mathbf{k}}\mathrm{e}^{i\phi}\\ d^{\prime}_{z}(\mathbf{k})\mathrm{e}^{-i\eta_{\mathbf{k}}}&[d^{\prime}_{x}(\mathbf{k})-id^{\prime}_{y}(\mathbf{k})]\mathrm{e}^{-i\eta_{\mathbf{k}}}&-\mu_{B}|\mathbf{H}|\sin\theta\mathrm{e}^{-i\phi}&-\xi_{\mathbf{k}}+\mu_{B}|\mathbf{H}|\cos\theta\end{pmatrix},$ (78)
@@
$\begin{split}\mathbf{d}(\mathbf{k})&=\mathbf{d}^{\prime}(\mathbf{k})\mathrm{e}^{i\eta_{\mathbf{k}}},\\ d_{x}(\mathbf{k})&=d_{x}^{\prime}(\mathbf{k})\mathrm{e}^{i\eta_{\mathbf{k}}},\\ d_{y}(\mathbf{k})&=d_{y}^{\prime}(\mathbf{k})\mathrm{e}^{i\eta_{\mathbf{k}}},\\ d_{z}(\mathbf{k})&=d_{z}^{\prime}(\mathbf{k})\mathrm{e}^{i\eta_{\mathbf{k}}}.\end{split}$ (79)
@@
$=$ $\begin{pmatrix}\cos{\frac{\theta}{2}}&\sin{\frac{\theta}{2}}&0&0\\ \sin{\frac{\theta}{2}}\mathrm{e}^{i\phi}&-\cos{\frac{\theta}{2}}\mathrm{e}^{i\phi}&0&0\\ 0&0&\cos{\frac{\theta}{2}}&\sin{\frac{\theta}{2}}\\ 0&0&\sin{\frac{\theta}{2}}\mathrm{e}^{-i\phi}&-\cos{\frac{\theta}{2}}\mathrm{e}^{-i\phi}\end{pmatrix}\begin{pmatrix}c_{\mathbf{k}+}\\ c_{\mathbf{k}-}\\ c_{-\mathbf{k}+}^{\dagger}\\ c_{-\mathbf{k}-}^{\dagger}\end{pmatrix},$
@@
$H^{\prime}_{\mathbf{k}}=\begin{pmatrix}\xi_{\mathbf{k}}+\mu_{B}|\mathbf{H}|&0&[-d_{1}(\mathbf{k})+id_{2}(\mathbf{k})]\mathrm{e}^{i\gamma_{\mathbf{k}}}&d_{3}(\mathbf{k})\mathrm{e}^{i\gamma_{\mathbf{k}}}\\ 0&\xi_{\mathbf{k}}-\mu_{B}|\mathbf{H}|&d_{3}(\mathbf{k})\mathrm{e}^{i\gamma_{\mathbf{k}}}&[d_{1}(\mathbf{k})+id_{2}(\mathbf{k})]\mathrm{e}^{i\gamma_{\mathbf{k}}}\\ [-d_{1}(\mathbf{k})-id_{2}(\mathbf{k})]\mathrm{e}^{-i\gamma_{\mathbf{k}}}&d_{3}(\mathbf{k})\mathrm{e}^{-i\gamma_{\mathbf{k}}}&-\xi_{\mathbf{k}}-\mu_{B}|\mathbf{H}|&0\\ d_{3}(\mathbf{k})\mathrm{e}^{-i\gamma_{\mathbf{k}}}&[d_{1}(\mathbf{k})-id_{2}(\mathbf{k})]\mathrm{e}^{-i\gamma_{\mathbf{k}}}&0&-\xi_{\mathbf{k}}+\mu_{B}|\mathbf{H}|\end{pmatrix},$ (81)